OpenStudy (123ab456c):

Solve for b. 8b = 48 A. 384 B. 56 C. 40 D. 6

OpenStudy (123ab456c):

can someone help me i will medal and fan

OpenStudy (alexandervonhumboldt2):

divide both sides by 8

OpenStudy (arihangdu):

6

OpenStudy (alexandervonhumboldt2):

8b/8=48/8 b=48/8=?

OpenStudy (123ab456c):

so the answer is 6

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

yes

OpenStudy (123ab456c):

thank you!!!

OpenStudy (arihangdu):

B=48/8=6